write an equation of a parabola that opens to the left has a vertex at the origin and a focus at (-9,0)
Answer by lwsshak3(11628)

write an equation of a parabola that opens to the left has a vertex at the origin and a focus at (-9,0)
**
Parabola as described is an equation of the standard form: (y-h)^2=-4p(x-h), (h,k)=(x,y) coordinates of the vertex.
For given parabola:
vertex:(0,0)
axis of symmetry: x-axis or y=0
p=9 (distance from focus to vertex on axis of symmetry)
4p=36
equation:
y^2=-36x